The speed of a particle changes from √5m/s to 2√5m/s in a time 't' . If the magnitude of change in its velocity is 5m/s the angle between the initial and final velocities of the particle is
Ans:90°
1
Expert's answer
2020-03-13T10:56:24-0400

Initial speed"=\sqrt{5} m/sec"

Final speed "=2\sqrt{5} m/sec"

And magnitude of change in velocity is "5m\/sec" which is equal to "\sqrt{(\sqrt{5})^2+({2\sqrt{5}})^2}" which means that magnitude in change of velocity is the unique case with velocities at different states do not have components in a common direction.

For example,Initial velocity"=\sqrt{5}\hat {i}"

Final velocity"=2\sqrt{5} \hat{j}"

Magnitude of change in velocity"=|2\sqrt{5}\hat{j}-\sqrt{5}\hat{i}|=\sqrt{(\sqrt{5})^2+({2\sqrt{5}})^2}=5m/sec"

Dot product of initial and final velocities"=(\sqrt{5}\hat{i}).(2\sqrt{5}\hat{j})=0"

or we can say the velocity at different states are perpendicular to each other or angle between them is "90\degree."